NPP Projection Report 3/15 Open
To do this evening/ was completed: Ordered us a new external hard drive since DCPIII has officially kicked the bucket.
Presentation Issues:
Preshow Clips/Playlist: Transferred the complete They Cloned Tyrone playlist from T2 to the LMS, along with the feature. Reminder to please ingest all content to the LMS (unless time simply doesn't permit) and build playlists on there, just simplifies things.
Scheduled Snow White and Alto Knights for Thursday.
Swapped most playlists on Thurs 3/20 for the 0321 versions, since just about every 0314 version has the Alto Knights or Snow White trailer.
Built REC and Conquest playlists.
Trailer Updates:
Special Events/ AV set up:
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Ripped the Watership Down blu-ray and started converting it to DCP. This is going onto the Fugazi hard drive (not ours, though we've had it for over a year now). Since we are currently without a portable external which is NTFS formatted, you'll have to copy the finished DCP elsewhere to ingest to the LMS, I'd recommend the "smb" share folder.
Outgoing/Shipping (content): Forgot to note in yesterday's blog that Pariah was picked up by DHL
Print/video Inspection:
Asset tracking doc updated/checked?:
Maintenance/cleaning: Put a new HDMI cable in T4, the old one wasn't working. Tested it quickly with the blu-ray player b/w rounds and then ran it along the wall to the rack. Will tackle T7's next week, the new long cable for it is here.
Trailer Updates:
Special Events/ AV set up:
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Ripped the Watership Down blu-ray and started converting it to DCP. This is going onto the Fugazi hard drive (not ours, though we've had it for over a year now). Since we are currently without a portable external which is NTFS formatted, you'll have to copy the finished DCP elsewhere to ingest to the LMS, I'd recommend the "smb" share folder.
Outgoing/Shipping (content): Forgot to note in yesterday's blog that Pariah was picked up by DHL
Print/video Inspection:
Asset tracking doc updated/checked?:
Maintenance/cleaning: Put a new HDMI cable in T4, the old one wasn't working. Tested it quickly with the blu-ray player b/w rounds and then ran it along the wall to the rack. Will tackle T7's next week, the new long cable for it is here.
Cue sheets updated/checked (every shift): Checked over tomorrow, sheet & schedule. We're all set.
Made the Snow White cue sheet formula and put those shows on Thursday's cue sheet.
Added the intro and discussion timing to They Clone Tyrone on Thursday's sheet.
Filled in credits for Gremlins 2 and Uptown Girls
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment